# Umlaut information

### v1

@@ IV.2 Zero-rate limit of the sphere-packing bound
IV.3 An explicit formula and additivity
Convention 18.
According to the usual conventions $00\log 0=$ 0 and $\exp(-\infty)=0$, we introduce the following notation. Let $\pazocal{W}$ be a discrete memoryless channel from $\mathcal{X}$ to $\mathcal{Y}$ and let $P_{X}\in\mathcal{P}(\mathcal{X})$. Then we define
$\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right)\coloneqq\sum_{y\in\mathcal{Y}_{\pazocal{W},P_{X}}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}P_{X}(x)\log\pazocal{W}(y|x)\right),$ (89)
Proposition 19.
$U(\pazocal{W})=-\log\min_{P_{X}}\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right)$ (90)
according to Convention 18, or, equivalently,
$\exp\left[-U(\pazocal{W})\right]$ $=\min_{P_{X}}\sum_{y\in\mathcal{Y}}\prod_{x\in\mathcal{X}}\pazocal{W}(y|x)^{P_{X}(x)},$ (91)
where we set $10^{0}=$ 1. It also holds that
$U(\pazocal{W})=\min_{Q_{Y}}\max_{x\in\mathcal{X}}D\left(Q_{Y}\middle\|\pazocal{W}(\,\cdot\,|x)\right)=\min_{Q_{Y}}\max_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}Q_{Y}(y)\log\frac{Q_{Y}(y)}{\pazocal{W}(y|x)}.$ (92)
Before proving the proposition, we need to state a minimax lemma.
Lemma 20 ([22, Theorem 5.2]).
Let $A$ be a compact, convex subset of a Hausdorff topological vector space $U$, and let $B$ be a convex subset of the linear space $V$. Let $f:A\times B\to(-\infty,+\infty]$ be lower semicontinuous on $A$ for all fixed $y\in B$, convex in the first variable, and concave in the second. Then
$\sup_{x\in A}\inf_{y\in B}f(x,y)=\inf_{y\in B}\sup_{x\in A}f(x,y).$ (93)
Proof of Proposition 19.
$U(\pazocal{W})$ $=\max_{P_{X}}\left(-H(P_{X})-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\left(P_{X}(x)\pazocal{W}(y|x)\right)\right)\right)$ (94)
$=\max_{P_{X}}-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right)$
$U(\pazocal{W})$ $=\max_{P_{X}}\min_{Q_{Y}}D(P_{X}Q_{Y}\|P_{XY})$ (95)
$=\max_{P_{X}}\min_{Q_{Y}}\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}(x)Q_{Y}(y)\log\frac{P_{X}(x)Q_{Y}(y)}{P_{X}(x)\pazocal{W}(y|x)}$
The expression to be optimised is linear in $P_{X}$ and convex in $Q_{Y}$, since it is the sum of two convex functions. Furthermore, $Q_{Y}\mapsto D(P_{X}Q_{Y}\|P_{XY})$ is lower semicontinuous [23, Theorem 15]. We can therefore apply Lemma 20 to get
$U(\pazocal{W})$ $=\min_{Q_{Y}}\max_{P_{X}}\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}(x)Q_{Y}(y)\log\frac{Q_{Y}(y)}{\pazocal{W}(y|x)}$ (96)
which concludes the proof. ∎
Corollary 21 ((Additivity of the channel umlaut information)).
$\big{(}\pazocal{W}_{1}\times\pazocal{W}_{2})(y_{1},y_{2}|x_{1},x_{2})\coloneqq\pazocal{W}_{1}(y_{1}|x_{1})\pazocal{W}_{2}(y_{2}|x_{2})$ (97)
$U(\pazocal{W}_{1}\times\pazocal{W}_{2})=U(\pazocal{W}_{1})+U(\pazocal{W}_{2}).$ (98)
Proof.
By (90), we immediately see that
$U(\pazocal{W}_{1}\times\pazocal{W}_{2})$ $=-\min_{P_{X_{1}X_{2}}}\log\sum_{\begin{subarray}{c}y_{1}\in\mathcal{Y}_{1}\\ y_{2}\in\mathcal{Y}_{2}\end{subarray}}\exp\left(\sum_{\begin{subarray}{c}x_{1}\in\mathcal{X}_{1}\\ x_{2}\in\mathcal{X}_{2}\end{subarray}}P_{X_{1}X_{2}}(x_{1},x_{2})\left(\log\pazocal{W}_{1}(y_{1}|x_{1})+\log\pazocal{W}_{2}(y_{2}|x_{2})\right)\right)$ (99)
=$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}-\min_{P_{X_{1}}}\log\sum_{y_{1}\in\mathcal{Y}_{1}}\exp\left(\sum_{x_{1}\in\mathcal{X}_{1}}P_{X_{1}}(x_{1})\log\pazocal{W}_{1}(y_{1}|x_{1})\right)$
$=U(\pazocal{W}_{1})+U(\pazocal{W}_{2})$
IV.4 Operational interpretation in non-signalling–assisted communication
@@
Consider a source of messages represented by a random variable $W$ taking values in an alphabet =$\mathcal{M}=\{1,\dots,M\}$ of size $M$. A code $\pazocal{Z}$ is defined by a joint probability distribution
$\pazocal{Z}(x,\hat{w}|y,w)=\mathbb{P}\left[X=x,\,\hat{W}=\hat{w}\,|\,Y=y,\,W=w\right],$ (100)
@@
    $\mathbb{P}\left[\hat{W}=\hat{w}\,|\,W=w,\,Y=y\right]$ $=\mathbb{P}\left[\hat{W}=\hat{w}\,|\,Y=y\right],$ (101)
@@
$\varepsilon^{\Omega}(M,\pazocal{W})\coloneqq\min_{\pazocal{Z}\in\{\Omega\text{-assisted codes}\}}\left\{\mathbb{P}\left[W\neq\hat{W}\right]\quad\text{ with }\quad Y|X\sim\pazocal{W},\quad X\hat{W}|YW\sim\pazocal{Z}\right\}$ (102)
@@
$M^{\Omega}(\varepsilon,\pazocal{W})\coloneqq\max_{\pazocal{Z}\in\{\Omega\text{-assisted codes}\}}\left\{M\,:\,\mathbb{P}[\hat{W}\neq W]\leq\varepsilon\quad\text{ with }\quad Y|X\sim\pazocal{W},\quad X\hat{W}|YW\sim\pazocal{Z}\right\}.$ (103)
@@
$\varepsilon^{\rm NS}(M,\pazocal{W})\leq\varepsilon^{\emptyset}(M,\pazocal{W}),\qquad M^{\emptyset}(\varepsilon,\pazocal{W})\leq M^{\rm NS}(\varepsilon,\pazocal{W}).$ (104)
@@
Definition 24 (($\Omega$-assisted zero-rate error exponent)).
@@
$E^{\Omega}(r,\pazocal{W})\coloneqq\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon^{\Omega}(\,\exp({rn}),\pazocal{W}^{\times n})\,.$ (105)
@@
$E^{\Omega}(0^{+},\pazocal{W})$ $\coloneqq\liminf_{r\to 0^{+}}E^{\Omega}(r,\pazocal{W})\,.$ (106)
@@
$E_{0}^{\Omega}(\pazocal{W})\coloneqq\liminf_{M\to\infty}\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon^{\Omega}(M,\pazocal{W}^{\times n}),$ (107)
which is to be understood as the error exponent with constant, but arbitrarily large message size $M\to\infty$ (similar to what is considered in [25, Equation (1)]). Note that for any fixed $M\geq 1$ and $r>0$,
$\varepsilon^{\Omega}(\,\exp(rn),\pazocal{W}^{\times n})\geq\varepsilon^{\Omega}(\,M,\pazocal{W}^{\times n})\qquad\forall\ n\geq\frac{\log M}{r}.$ (108)
@@
$\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon^{\Omega}(\exp({rn}),\pazocal{W}^{\times n})\leq\liminf_{n\to\infty}-\frac{1}{n}\varepsilon^{\Omega}(\,M,\pazocal{W}^{\times n}),$ (109)
@@
$E^{\Omega}(0^{+},\pazocal{W})\leq E_{0}^{\Omega}(\pazocal{W}).$ (110)
Further, it follows from [26, Equation (1.55)] that in fact $E^{\emptyset}(0^{+},\pazocal{W})=E^{\emptyset}_{0}(\pazocal{W})$, and we will prove as part of the forthcoming Theorem 26 that also $E^{\mathrm{\rm NS}}(0^{+},\pazocal{W})=E^{\mathrm{\rm NS}}_{0}(\pazocal{W})$. As such, $E^{\Omega}(0^{+},\pazocal{W})$ and $E_{0}^{\Omega}(\pazocal{W})$ are identical quantifiers for the zero-rate error exponents considered in this work.
@@
$E_{0}^{\rm NS}(\pazocal{W})$ $=\lim_{\delta\to 0}\liminf_{n\to\infty}\frac{1}{n}\max_{P_{X^{n}}}\min_{Q_{Y^{n}}}D^{\delta}_{H}(P_{X^{n}}Q_{Y^{n}}\|P_{X^{n}Y^{n}})$ (112)
@@
This essentially follows from the considerations in [8] and the one-shot result has been previously stated as [27, Proposition B.1]. We give a self-contained proof in our notation in Appendix A. The following theorem is our main result in this section.
@@
Given a discrete, memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$, it holds that $E_{0}^{\rm NS}(\pazocal{W})=E^{\rm NS}(0^{+},\pazocal{W})=E_{\mathrm{sp}}(0^{+},\pazocal{W})=U(\pazocal{W}).$ (113)
@@
The last equality $E_{\mathrm{sp}}(0^{+},\pazocal{W})=U(\pazocal{W})$ was shown in Proposition 17. The second equality $E^{\rm NS}(0^{+},\pazocal{W})=E_{\mathrm{sp}}(0^{+},\pazocal{W})$ can then also be deduced from taking the $r\to 0$ limit of $E^{\mathrm{NS}}(r,\pazocal{W})=E_{\mathrm{sp}}(r,\pazocal{W})$, which is known from [28, Theorem 24] (also see the discussion in [29, Theorem 4.1]). We will show that $E_{0}^{\rm NS}(\pazocal{W})$ and $E^{\rm NS}(0^{+},\pazocal{W})$ are both equal to $U(\pazocal{W})$.
@@
$P^{\times n}_{X}(x_{1},\dots,x_{n})=P_{X}(x_{1})\cdots P_{X}(x_{n})$ (114)
@@
$E^{\rm NS}(0^{+},\pazocal{W})$ $=\lim_{r\to 0}\liminf_{n\to\infty}\frac{1}{n}\max_{P_{X^{n}}}\min_{Q_{Y^{n}}}D^{\exp(-rn)}_{H}(P_{X}^{\times n}Q_{Y^{n}}\|P_{X^{n}Y^{n}})$ (115)
@@
=$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{=}}\liminf_{n\to\infty}\frac{1}{n}U_{\alpha}(X^{n};Y^{n})$
=$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iii)}}}}{{=}}U_{\alpha}(X;Y)$
@@
$D^{\varepsilon}_{H}(P\|Q)\geq D_{\alpha}(P\|Q)+\frac{\alpha}{1-\alpha}\log\frac{1}{\varepsilon}$ (116)
@@
$E^{\rm NS}(0^{+},\pazocal{W})$ $\geq\max_{P_{X}}\limsup_{\alpha<1}U_{\alpha}(X;Y)\,\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iv)}}}}{{=}}\,\max_{P_{X}}U(X;Y)=U(\pazocal{W})$ (117)
@@
$E_{0}^{\rm NS}(\pazocal{W})$ $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(v)}}}}{{\leq}}\lim_{\delta\to 0}\liminf_{n\to\infty}\max_{P_{X^{n}}}\min_{Q_{Y^{n}}}\frac{1}{n}\frac{1}{1-\delta}\left(1+D(P_{X^{n}}Q_{Y^{n}}\|P_{X^{n}Y^{n}})\right)$ (118)
@@
=$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(vi)}}}}{{=}}U(\pazocal{W}),$
where in (v) we have used the upper bound (see e.g. [30, Equation (2.251)]).
$D_{H}^{\varepsilon}(P\|Q)\leq\frac{1}{1-\varepsilon}(1+D(P\|Q))$ (119)
@@
$U(\pazocal{W})\leq E^{\rm NS}(0^{+},\pazocal{W})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(vii)}}}}{{\leq}}E_{0}^{\rm NS}(\pazocal{W})\leq U(\pazocal{W}),$ (120)
@@
IV.5 Operational interpretation in unassisted communication

### v2

@@ IV.4 Zero-rate limit of the sphere-packing bound
IV.5 Operational interpretation in non-signalling–assisted communication
@@
Consider a source of messages represented by a random variable $W$ taking values in an alphabet $\mathcal{M}=\{1,\dots,M\}$ of size $M$. A code $\pazocal{Z}$ is defined by a joint probability distribution
$\hskip 0.0pt\pazocal{Z}(x,\hat{w}|y,w)=\mathbb{P}\left[X=x,\,\hat{W}=\hat{w}\,|\,Y=y,\,W=w\right],$ (131)
@@
    $\hskip 0.0pt\mathbb{P}\left[\hat{W}=\hat{w}\,|\,W=w,\,Y=y\right]$ $=\mathbb{P}\left[\hat{W}=\hat{w}\,|\,Y=y\right],$ (132)
@@
$\hskip 0.0pt\varepsilon^{\Omega}(M,\pazocal{W})\coloneqq\min_{\pazocal{Z}\in\{\Omega\text{-assisted codes}\}}\left\{\mathbb{P}\left[W\neq\hat{W}\right]\quad\text{ with }\quad Y|X\sim\pazocal{W},\quad X\hat{W}|YW\sim\pazocal{Z}\right\}$ (133)
@@
$\hskip 0.0ptM^{\Omega}(\varepsilon,\pazocal{W})\coloneqq\max_{\pazocal{Z}\in\{\Omega\text{-assisted codes}\}}\left\{M\,:\,\mathbb{P}[\hat{W}\neq W]\leq\varepsilon\quad\text{ with }\quad Y|X\sim\pazocal{W},\quad X\hat{W}|YW\sim\pazocal{Z}\right\}.$ (134)
@@
$\hskip 0.0pt\varepsilon^{\rm NS}(M,\pazocal{W})\leq\varepsilon^{\emptyset}(M,\pazocal{W}),\qquad M^{\emptyset}(\varepsilon,\pazocal{W})\leq M^{\rm NS}(\varepsilon,\pazocal{W}).$ (135)
@@
Definition 28 $($$\Omega$-assisted zero-rate error exponent)).
@@
$\hskip 0.0ptE^{\Omega}(r,\pazocal{W})\coloneqq\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon^{\Omega}(\,\exp({rn}),\pazocal{W}^{\times n})\,.$ (136)
@@
$\hskip 0.0ptE^{\Omega}(0^{+},\pazocal{W})$ $\coloneqq\liminf_{r\to 0^{+}}E^{\Omega}(r,\pazocal{W})\,.$ (137)
@@
$\hskip 0.0ptE_{0}^{\Omega}(\pazocal{W})\coloneqq\liminf_{M\to\infty}\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon^{\Omega}(M,\pazocal{W}^{\times n}),$ (138)
which is to be understood as the error exponent with constant, but arbitrarily large message size $M\to\infty$ (similar to what is considered in [27, Equation (1)]). We will shortly see, in fact, that the asymptotic results do not depend on the message size $M$ and hence we do not need to take $M$ going to infinity: it will suffice to fix an arbitrary $M\geq 2$. Note that for any fixed $M\geq 1$ and $r>0$,
$\hskip 0.0pt\varepsilon^{\Omega}(\,\exp(rn),\pazocal{W}^{\times n})\geq\varepsilon^{\Omega}(\,M,\pazocal{W}^{\times n})\qquad\forall\ n\geq\frac{\log M}{r}.$ (139)
@@
$\hskip 0.0pt\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon^{\Omega}(\exp({rn}),\pazocal{W}^{\times n})\leq\liminf_{n\to\infty}-\frac{1}{n}\varepsilon^{\Omega}(\,M,\pazocal{W}^{\times n}),$ (140)
@@
$\hskip 0.0ptE^{\Omega}(0^{+},\pazocal{W})\leq E_{0}^{\Omega}(\pazocal{W}).$ (141)
Further, it follows from [28, Equation (1.55)] that in fact $E^{\emptyset}(0^{+},\pazocal{W})=E^{\emptyset}_{0}(\pazocal{W})$, and we will prove as part of the forthcoming Theorem 30 that also $E^{\mathrm{\rm NS}}(0^{+},\pazocal{W})=E^{\mathrm{\rm NS}}_{0}(\pazocal{W})$. As such, $E^{\Omega}(0^{+},\pazocal{W})$ and $E_{0}^{\Omega}(\pazocal{W})$ are identical quantifiers for the zero-rate error exponents considered in this work.
@@
$\hskip 0.0ptE_{0}^{\rm NS}(\pazocal{W})$ $=\lim_{\delta\to 0}\liminf_{n\to\infty}\frac{1}{n}\max_{P_{X^{n}}}\min_{Q_{Y^{n}}}D^{\delta}_{H}(P_{X^{n}}Q_{Y^{n}}\|P_{X^{n}Y^{n}})$ (143)
@@
This essentially follows from the considerations in [8] and the one-shot result has been previously stated as [29, Proposition B.1]. We give a self-contained proof in our notation in Appendix A. The following theorem is our main result in this section.
@@
Given a discrete, memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$, it holds that $\hskip 0.0ptE_{0}^{\rm NS}(\pazocal{W})=E^{\rm NS}(0^{+},\pazocal{W})=E_{\mathrm{sp}}(0^{+},\pazocal{W})=U(\pazocal{W}).$ (144)
@@
The last equality $E_{\mathrm{sp}}(0^{+},\pazocal{W})=U(\pazocal{W})$ was shown in Proposition 25. The second equality $E^{\rm NS}(0^{+},\pazocal{W})=E_{\mathrm{sp}}(0^{+},\pazocal{W})$ can then also be deduced from taking the $r\to 0$ limit of $E^{\mathrm{NS}}(r,\pazocal{W})=E_{\mathrm{sp}}(r,\pazocal{W})$, which is known from [19, Theorem 24] (also see the discussion in [30, Theorem 4.1]). We will show that $E_{0}^{\rm NS}(\pazocal{W})$ and $E^{\rm NS}(0^{+},\pazocal{W})$ are both equal to $U(\pazocal{W})$.
@@
$\hskip 0.0ptP^{\times n}_{X}(x_{1},\dots,x_{n})=P_{X}(x_{1})\cdots P_{X}(x_{n})$ (145)
@@
$\hskip 0.0ptE^{\rm NS}(0^{+},\pazocal{W})$ $=\lim_{r\to 0}\liminf_{n\to\infty}\frac{1}{n}\max_{P_{X^{n}}}\min_{Q_{Y^{n}}}D^{\exp(-rn)}_{H}(P_{X}^{\times n}Q_{Y^{n}}\|P_{X^{n}Y^{n}})$ (146)
@@
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{=}}\liminf_{n\to\infty}\frac{1}{n}U_{\alpha}(X^{n};Y^{n})$
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iii)}}}}{{=}}U_{\alpha}(X;Y)$
@@
$\hskip 0.0ptD^{\varepsilon}_{H}(P\|Q)\geq D_{\alpha}(P\|Q)+\frac{\alpha}{1-\alpha}\log\frac{1}{\varepsilon}$ (147)
@@
$\hskip 0.0ptE^{\rm NS}(0^{+},\pazocal{W})$ $\geq\max_{P_{X}}\limsup_{\alpha<1}U_{\alpha}(X;Y)\,\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iv)}}}}{{=}}\,\max_{P_{X}}U(X;Y)=U(\pazocal{W})$ (148)
@@
$\hskip 0.0ptE_{0}^{\rm NS}(\pazocal{W})$ $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(v)}}}}{{\leq}}\lim_{\delta\to 0}\liminf_{n\to\infty}\max_{P_{X^{n}}}\min_{Q_{Y^{n}}}\frac{1}{n}\frac{1}{1-\delta}\left(1+D(P_{X^{n}}Q_{Y^{n}}\|P_{X^{n}Y^{n}})\right)$ (149)
@@
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(vi)}}}}{{=}}U(\pazocal{W}),$
where in (v) we have again used the achievability of the meta-converse in Proposition 29 as well as the upper bound (see e.g. [31, Equation (2.251)])
$\hskip 0.0ptD_{H}^{\varepsilon}(P\|Q)\leq\frac{1}{1-\varepsilon}(1+D(P\|Q)),$ (150)
@@
$\hskip 0.0ptU(\pazocal{W})\leq E^{\rm NS}(0^{+},\pazocal{W})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(vii)}}}}{{\leq}}E_{0}^{\rm NS}(\pazocal{W})\leq U(\pazocal{W}),$ (151)
@@
A ‘strong converse’ extension of the operational interpretation established in Theorem 30 can be formulated as follows. For a fixed message size $M$, let
$\hskip 0.0ptE_{0}^{\Omega}(\pazocal{W},M)\coloneqq\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon^{\Omega}(M,\pazocal{W}^{\times n}).$ (152)
By definition of $E_{0}^{\Omega}(\pazocal{W})$ in (138), this means that
$\hskip 0.0ptE_{0}^{\Omega}(\pazocal{W})=\liminf_{M\to\infty}E_{0}^{\Omega}(\pazocal{W},M).$ (153)
Then, we can prove the following statement.
Theorem 31 ((Strong converse for the non-signalling–assisted zero-rate error exponent)).
Given a discrete, memoryless channel $\pazocal{W}$ from $\mathcal{X}$ to $\mathcal{Y}$, for any fixed $M\geq 2$ it holds that
$\hskip 0.0ptE_{0}^{\rm NS}(\pazocal{W},M)=U(\pazocal{W}).$ (154)
Proof.
By the very definition of $E_{0}^{\rm NS}(\pazocal{W},M)$, we have that
$\hskip 0.0ptE_{0}^{\rm NS}(\pazocal{W},M)\geq E_{0}^{\rm NS}(\pazocal{W})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}U(\pazocal{W}),$ (155)
where in (i) we have used Theorem 30. Using the achievability of the meta-converse (Proposition 29), we now see that
$\hskip 0.0pt-\log\varepsilon^{\rm NS}(M,\pazocal{W})=\max_{P_{X}}\min_{Q_{Y}}D^{1/M}_{H}(P_{X}Q_{Y}\|P_{XY}),$ (156)
so that
$\hskip 0.0ptE_{0}^{\rm NS}(\pazocal{W},M)=\liminf_{n\to\infty}\max_{P_{X^{n}}}\min_{Q_{Y^{n}}}\frac{1}{n}D^{1/M}_{H}(P_{X^{n}}Q_{Y^{n}}\|P_{X^{n}Y^{n}}).$ (157)
Hence
$\hskip 0.0ptE_{0}^{\rm NS}(\pazocal{W},M)$ $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{\leq}}\liminf_{n\to\infty}\max_{P_{X^{n}}}\min_{Q_{Y^{n}}}\frac{1}{n}\left(D_{\alpha}(P_{X^{n}}Q_{Y^{n}}\|P_{X^{n}Y^{n}})+\frac{\alpha}{1-\alpha}\log\frac{1}{1-1/M}\right)$ (158)
$=\liminf_{n\to\infty}\frac{1}{n}U_{\alpha}(\pazocal{W}^{\times n})$
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iii)}}}}{{=}}U_{\alpha}(\pazocal{W}),$
where in (ii) we have used the upper bound (see e.g. [32])
$\hskip 0.0ptD_{H}^{\varepsilon}(p\|q)\leq D_{\alpha}(p\|q)+\frac{\alpha}{\alpha-1}\log\frac{1}{1-\varepsilon},$ (159)
which holds for $0<\varepsilon<1$ and $\alpha\in(1,\infty)$, and in (iii) we used the additivity of the Rényi $\alpha$-umlaut information of a channel (Corollary 23). By arbitrariness of $\alpha>1$, we take the limit
$\hskip 0.0ptE_{0}^{\rm NS}(\pazocal{W},M)$ $\leq\lim_{\alpha\to 1^{+}}U_{\alpha}(\pazocal{W})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iv)}}}}{{=}}U(\pazocal{W}),$ (160)
where in (iv) we have used Corollary 24. This concludes the proof. ∎
IV.6 Operational interpretation in unassisted communication